OpenStudy (mattyboyy):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? A) (x + 2)^2 + (y +3)^2 = 9 B) (x - 2)^2 + (y - 3)^2 = 16 C) (x - 2)^2 + (y - 3)^2 = 4 D) (x + 2)^2 + (y + 3)^2 = 16 E) (x- 2)^2 + (y +3)^2 = 9

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2

OpenStudy (anonymous):

the centre (a,b)=(-2,-3)

OpenStudy (anonymous):

(x-(-2))^2+(y-(-3))^2=r^2 (x+2)^2+(y+3)^2=r^2

OpenStudy (anonymous):

r^2=(x-a)^2+(y-b)^2 =(-2-(-2))^2+(0-(-3))^2 =(0)^2+(3)^2=9 (x+2)^2+(y+3)^2=9

OpenStudy (anonymous):

A is the answer